# D2USt3R: Enhancing 3D Reconstruction with 4D Pointmaps for Dynamic Scenes

### v1

@@ Abstract
We address the task of 3D reconstruction in dynamic scenes, where object motions degrade the quality of previous 3D pointmap regression methods, such as DUSt3R, originally designed for static 3D scene reconstruction. Although these methods provide an elegant and powerful solution in static settings, they struggle in the presence of dynamic motions that disrupt alignment based solely on camera poses. To overcome this, we propose D2USt3R that regresses 4D pointmaps that simultaneiously capture both static and dynamic 3D scene geometry in a feed-forward manner. By explicitly incorporating both spatial and temporal aspects, our approach successfully encapsulates spatio-temporal dense correspondence to the proposed 4D pointmaps, enhancing downstream tasks. Extensive experimental evaluations demonstrate that our proposed approach consistently achieves superior reconstruction performance across various datasets featuring complex motions.
Figure 1: Teaser. Given a pair of input views, our D2USt3R  accurately establishes dense correspondence not only in static regions but also in dynamic regions, enabling full reconstruction of a dynamic scene via our proposed 4D pointmap. As highlighted by red dots $\bullet$ in the pointmap, DUSt3R [42] and MonST3R [48] align pointmaps solely based on camera motion, causing corresponding 2D pixels to become misaligned in 3D space. We compare the cross-attention maps, established correspondence fields, and estimated depth maps produced by our D2USt3R against baseline methods.
Co-first authors.Co-corresponding authors.
@@ 1 Introduction
Recovering 3D scene geometry from images remains a central challenge in computer vision. Traditional approaches, such as Structure-from-Motion (SfM) [32] and Multi-View Stereo (MVS) [34], have achieved impressive results in this context. While these methods are originally designed for recovering precise 3D scene geometry, they often struggle with scenes that include dynamic objects, symmetries, imagery with minimal overlapping and textureless regions [25, 33, 3, 9].
@@
Recent approaches [42, 21, 40] leverage deep learning to streamline the 3D reconstruction pipeline and enhance robustness. DUSt3R [42], as a pioneering method, introduced a unified learning-based framework for dense stereo 3D reconstruction by jointly addressing image matching, essential matrix estimation, and triangulation. Specifically, DUSt3R directly regresses 3D pointmaps that encode scene geometry, pixel-to-scene correspondences, and inter-view relationships, mitigating error accumulation typical in multi-stage pipelines. Despite its strengths in static scenarios, DUSt3R significantly struggles with dynamic scenes due to its rigidity assumption, as exemplified in Figure 1.
@@
Dynamic scenes, prevalent in real-world scenarios, pose significant challenges in 3D scene reconstruction task, as object motion disrupts the camera pose-based alignment [42], causing misaligned correspondences and inaccurate depth estimates. Recent methods such as MonST3R [48] tried to address dynamic 3D scene reconstruction by training on dynamic scene video data collections, but still relied on rigidly warped pointmaps within a common coordinate frame. Consequently, these approaches suffer from compromised correspondence learning for dynamic objects, impairing depth accuracy and robust geometry recovery.
In this paper, we propose Dynamic Dense Stereo 3D Reconstruction (D2USt3R), a novel feed-forward framework that directly regresses 4D pointmaps, simultaneously accounting for both spatial structure and motion to enable more reliable 3D reconstruction of both static and dynamic scene regions. This is achieved by directly introducing and aiming to establish dense correspondence fields between dynamic objects into the training process, ensuring consistent alignment of pointmaps across both static and dynamic regions. Compared to MonST3R [48], our model effectively captures dense correspondences between frames and enhances depth accuracy for dynamic elements by treating correspondence and reconstruction in dynamic scenes as entangled components. Our approach maintains the original DUSt3R [42] architecture, simply introducing dynamic motion awareness through a novel training signal. Experimental results and visual comparisons clearly demonstrate the significant improvement in reconstruction accuracy provided by our method.
Our key contributions are summarized as follows:
  • We introduce a novel 3D reconstruction framework that explicitly models dynamic motion through dynamically aligned 4D pointmaps, enabling unified reconstruction of complex scenes.
@@
  • To compensate for the missing direct 3D correspondences between dynamic objects, we propose a 3D alignment loss that effectively handles occlusions and motions.
@@
Reconstruction Correspondence
Methods Static Dynamic Static Dynamic
DUSt3R [42]
MASt3R [21] $\triangle$
MonST3R [48]
D2USt3R(Ours)
Table 1: Comparison of our method with existing approaches. Our proposed method uniquely handles both reconstruction and matching tasks consistently across static and dynamic scenarios. Note that MASt3R [21] employs a matching head to match points; however, it operates under the assumption of a static scene.
@@ 2 Related Work
Figure 2: Cross-attention visualization of DUSt3R [42] on static and dynamic scenes. We present visualizations of the source image’s attention maps corresponding to the query point (highlighted in red on the target image), both at each individual layer and as an average across layers. Although DUSt3R [42] effectively captures geometric correspondences during 3D reconstruction, it struggles to establish correspondences in regions exhibiting dynamic motion. This issue stems from its training regime, which assumes that both frames are static and can be modeled solely with rigid camera motion, thereby neglecting the complexities introduced by dynamic scenes.
Figure 3: Comparison of cross-attention on dynamic scene. Although MonST3R [48] is trained on dynamic video, its training signal remains identical to that of DUSt3R [42]. As a result, it fails to achieve reliable alignment between frames, ultimately limiting its 3D reconstruction performance. In contrast, our D2USt3R successfully establishes correspondences between dynamic frames, and thus has a stronger ability to estimate 3D shapes from dynamic motions.
@@
Classical 3D reconstruction methods typically follow a multi-stage pipeline to recover scene geometry and camera parameters from a set of uncalibrated images. Prominent examples include Structure-from-Motion (SfM) [49], and Simultaneous Localization and Mapping (SLAM) [8]. SfM incrementally reconstructs sparse 3D points through feature matching and bundle adjustment where SLAM simultaneously estimates camera trajectories and builds sparse or semi-dense maps in real-time. Building upon SfM, ParticleSfM [50] incorporate particle-based trajectory modeling to track object motions, paving the way for reconstruction in dynamic scenes. Despite all of these approaches showing remarkable performance, they typically require long processing time and resources for scene-specific optimization and often struggle with error accumulations from multi-stage pipelines.
@@
Building on top of estabilished correspondence fields [4, 5, 14, 16, 15], various learning-based approaches [40, 24, 12, 36, 18] have been proposed to reconstruct static 3D scenes by learning strong 3D priors, representing the scene as point clouds [24, 13], meshes [12, 41], voxels [36, 6] and 3DGS [17]. Recently, DUSt3R [42] notably provides a unified, feed-forward pipeline for dense stereo matching, geometry estimation, and triangulation by directly regressing structured 3D pointmaps. Although DUSt3R significantly improves reconstruction quality and efficiency by reducing cumulative errors, DUSt3R is inherently designed for static scenes, limiting its effectiveness in scenarios involving dynamic components.
@@
Dynamic scene reconstruction introduces additional complexities due to non-rigid transformations occurring across frames. Similarly to static scene reconstruction, several recent approaches [23, 48, 27] have employed learning-based methods to tackle dynamic scene reconstruction. Among these, MonST3R [48] directly fine-tunes DUSt3R using dynamic videos, enabling a feed-forward approach to reconstruct dynamic scenes. Despite demonstrating strong overall performance, MonST3R retains DUSt3R’s per-frame training paradigm and lacks explicit mechanisms for linking corresponding points across frames in dynamic scenes. As demonstrated in Table 1, while DUSt3R is trained specifically to align corresponding 3D points, MonST3R does not effectively preserve this alignment during finetuning for dynamic content. Consequently, dynamic regions can suffer from inconsistent depth estimations, arising from inadequate temporal constraints that fail to capture the intricate motion patterns of objects. Our method addresses this limitation by augmenting the existing feed-forward framework with motion-aware training objectives, explicitly enforcing consistent 3D point correspondences over time.
3 Preliminary - DUSt3R
Given a pair of input images $I^{1},I^{2}\in\mathbb{R}^{W\times H\times 3}$, DUSt3R [42] predicts a pair of 3D pointmaps $X^{1,1},X^{2,1}\in\mathbb{R}^{W\times H\times 3}$ for both images, each expressed in the camera coordinate system of $I^{1}$. To train the network in a supervised manner, ground-truth pointmaps for each image are defined in the coordinate space of the first camera. Specifically, given the camera intrinsics matrix $K\in\mathbb{R}^{3\times 3}$, world-to-camera pose matrices $P_{n},P_{m}\in\mathbb{R}^{4\times 4}$ for images $n$ and $m$, and a ground-truth depth map $D\in\mathbb{R}^{W\times H}$, the ground-truth pointmap is computed as $X^{n,m}=P_{m}P_{n}^{-1}h(K^{-1}D)$, where $h:(x,y,z)\mapsto(x,y,z,1)$ represents the transformation to homogeneous coordinates.
@@ 3 Preliminary - DUSt3R
where $v\in\{1,2\}$ denotes the input views and $i\in\mathcal{D}^{v}$ denotes valid pixel positions. The scaling factors $z=\text{norm}(X^{1,1},X^{2,1})$ and $\bar{z}=\text{norm}(\bar{X}^{1,1},\bar{X}^{2,1})$ are computed using the normalization function:
@@ 4 Methodology
Figure 4: Construction of alignment loss defined at static and dynamic regions. We propose a pipeline for generating a 4D pointmap while explicitly addressing occlusions caused by dynamic regions. (a) From the input images, we obtain optical flow refined via cycle consistency checks and derive a dynamic mask $M_{\text{dyn}}$. To align image $I^{2}$ with image $I^{1}$, we utilize: 1) the camera pose to align static regions in $I^{2}$, and 2) optical flow to align dynamic regions. The alignment process is conducted specifically within regions corresponding to the colored pixels. Through this process, we construct a 4D pointmap, enabling alignment in 3D space for all corresponding 2D pixels.
@@ 4.1 Motivation and overview
As visualized in Figure 1, DUSt3R [42] learns to predict 3D pointmaps primarily defined in static regions, enabling accurate and robust 3D reconstruction through precise stereo correspondences. Motivated by ZeroCo’s [1] finding that the cross-attention maps of DUSt3R inherently encode geometric correspondence information, we visualize the cross-attention maps for both static and dynamic scenes in Figure 2. However, the underlying assumption that reliable pointmaps exist only in static regions limits its applicability to dynamic scenes. To compensate for, recent methods, such as MonST3R [48], have attempted to address dynamic scenes by extending DUSt3R’s framework to support 4D reconstruction through augmented training datasets that include dynamic scenarios. Nevertheless, this approach overlooks the explicit correspondence between dynamic objects, resulting in inaccurate depth estimations. This omission impairs the ability to leverage dynamic objects as anchor points, which could otherwise reinforce the spatial structure and improve depth estimation of nearby static regions, ultimately hindering overall 3D reconstruction performance. Figure 3 visualizes these issues through attention maps, highlighting noisy correspondences that degrade reconstruction accuracy.
Motivated by these limitations, we propose extending 3D pointmaps to 4D pointmaps to explicitly encode motion dynamics. Our method ensures accurate alignment not only in static regions but also among dynamic objects. In the following sections, we describe our proposed 4D pointmap representation and outline our training procedure.
4.2 Occlusion and dynamic masks preparation
We find that existing datasets often do not include 3D correspondences between dynamic objects, which poses additional challenges in learning them. To alleviate this, we first employ off-the-shelf optical flow [44] to find dense 2D correspondences across image pairs. Note that if the ground-truth optical flows are available in the dataset, we instead use them. For the scenarios where occlusions are more frequent due to large camera baselines, we perform forward-backward consistency check [38, 47] to additionally obtain occlusion masks. This process is defined as follows:
$p^{\prime}_{2}=p_{1}+\mathbf{f}(p_{1}),\quad p^{\prime}_{1}=p^{\prime}_{2}+\mathbf{b}(p^{\prime}_{2}),\quad$ (4)
${M}_{\mathrm{occ}}=|p^{\prime}_{1}-p_{1}|>t,$
@@ 4.2 Occlusion and dynamic masks preparation
While we empirically find that occlusion mask provides reliable dense correspondences for supervision, due to frequent large changes in camera and dynamic objects motions, we find it helpful to introduce a dynamic mask ${M}_{\mathrm{dyn}}$, to encourage stable learning process by informing the network the regions of dynamic objects within the images. Specifically, the dynamic mask ${M}_{\mathrm{dyn}}$ is computed by comparing optical flow $\mathbf{f}$ with the expected flow induced purely by camera motion $\mathbf{f}_{\text{cam}}$. Given the depth map $D$, intrinsics matrix $K$, relative rotation and translation $R,T$, and pixel coordinates $p$, we define:
$\mathbf{f}_{\text{cam}}$ $=\pi(DKRK^{-1}p+KT)-p,$ (5)
$\mathcal{\mathcal{missing}}{M}_{\mathrm{dyn}}$ $=[\|\mathbf{f}_{\text{cam}}-\mathbf{f}\|<\tau],$
where $\pi:(x,y,z)\mapsto(x/z,y/z)$ denotes the projection operation, and $\tau$ serves as the dynamic threshold.
Dataset Domain ##\## of frames ##\## of Scenes Dynamics Dynamic mask Optical flow Ratio
Blinkvision Outdoor [22] Outdoors 6k 23 Realistic 38.75%
Blinkvision Indoor [22] Indoors 6k 24 Realistic 23.75%
PointOdyssey [51] Indoors & Outdoors 200k 131 Realistic 12.5%
TartanAir [43] Indoors & Outdoors 1000k 163 None 12.5%
Spring [28] Outdoors 6k 37 Realistic 12.5%
Table 2: Training datasets utilized for fine-tuning the pretrained DUSt3R [42] on dynamic scenes. All datasets consists of synthetic scenes and provide both camera pose and depth information, with most containing dynamic objects. For the PointOdyssey dataset, we excluded scenes containing smoke or those with inaccurate segmentation masks. For the Blinkvision Outdoor dataset, scenes with incomplete dataset annotations were omitted. Details regarding each scene are documented in the supplementary material A. Training dataset.
4.3 4D pointmaps regression

### v2

# Enhancing 3D Reconstruction for Dynamic Scenes

@@ Abstract
In this work, we address the task of 3D reconstruction in dynamic scenes, where object motions frequently degrade the quality of previous 3D pointmap regression methods, such as DUSt3R, that are originally designed for static 3D scene reconstruction. Although these methods provide an elegant and powerful solution in static settings, they struggle in the presence of dynamic motions that disrupt alignment based solely on camera poses. To overcome this, we propose D2USt3R that directly regresses Static-Dynamic Aligned Pointmaps (SDAP) that simultaneiously capture both static and dynamic 3D scene geometry. By explicitly incorporating both spatial and temporal aspects, our approach successfully encapsulates 3D dense correspondence to the proposed pointmaps, enhancing downstream tasks. Extensive experimental evaluations demonstrate that our proposed approach consistently achieves superior 3D reconstruction performance across various datasets featuring complex motions.
Figure 1: Teaser. Given a pair of input views, D2USt3R accurately establishes dense correspondence not only in static regions but also in dynamic regions, and enables 3D reconstruction of a dynamic scene via our proposed static-dynamic aligned pointmap. The colored $\bullet$ pointmaps highlight that DUSt3R wang2024dust3r and MonST3R zhang2024monst3r align pointmaps solely based on camera motion, causing corresponding 2D pixels within dynamic object misaligned. We also compare the cross-attention maps, established correspondence fields, and estimated depth maps produced by D2USt3R against baseline methods wang2024dust3r ; zhang2024monst3r , where our method shows higher precision.
@@ 1 Introduction
Recovering 3D scene geometry from images remains a core problem in computer vision. Traditional approaches, such as Structure-from-Motion (SfM) schonberger2016structure and Multi-View Stereo (MVS) seitz2006comparison , have achieved impressive results in this context. While these methods are originally designed for recovering precise 3D scene geometry, they often struggle with scenes that include objects, symmetries, imagery with minimal overlapping and textureless regions lindenberger2021pixel ; schops2017multi ; campos2021orb ; furukawa2015multi .
@@
To overcome these limitations, recent approaches wang2024dust3r ; leroy2024grounding ; wang2024vggsfm leverage a learning framework to streamline the 3D reconstruction pipeline and enhance performance. DUSt3R wang2024dust3r , as a pioneering method, introduced a unified learning-based framework for dense stereo 3D reconstruction. Specifically, DUSt3R directly regresses 3D pointmaps that encode scene geometry, pixel-to-scene correspondences, and inter-view relationships, mitigating error accumulation typical in multi-stage pipelines. Despite its strengths in static scenarios, DUSt3R significantly struggles with dynamic scenes due to its rigidity assumption, as exemplified in Figure 1.
@@
Dynamic scenes, prevalent in real-world scenarios, pose significant challenges in 3D scene reconstruction task, as object motions disrupts the camera pose-based alignment wang2024dust3r , causing misaligned correspondences and inaccurate depth estimates in dynamic objects, which further degrades reconstruction accuracy in static regions. While recent methods such as MonST3R zhang2024monst3r extends training to dynamic‐scene video collections to account for dynamic objects, it still models all pointmaps as if generated by a single global rigid transformation. Consequently, these approaches suffer from compromised correspondence learning for dynamic objects, in turn impairing depth accuracy and robust geometry recovery.
In this paper, we propose Dynamic Dense Unconstrained Stereo 3D Reconstruction (D2USt3R), a novel feed-forward framework that directly regresses Static-Dynamic Aligned Pointmaps (SDAP), simultaneously accounting for both spatial structures and temporal motions to enable more reliable 3D reconstruction of both static and dynamic regions. Unlike MonST3R zhang2024monst3r , which overlooks correspondences on moving objects, our model captures dense inter-frame matches by treating correspondence and reconstruction as a unified problem in dynamic scenes. We achieve this with a novel training scheme that applies separate supervisory signals to static and dynamic regions, signals that are further stabilized and localized to regions of interests by our occlusion and dynamic masks. Experimental results and visual comparisons confirm that our approach delivers a significant boost in reconstruction accuracy.
Our contributions are summarized as follows:
  • Our approach, D2USt3R, captures dynamic motion by leveraging static–dynamic aligned pointmaps, allowing for the comprehensive 3D reconstruction of all scene elements in any environment.
@@
  • To compensate for the missing direct 3D correspondences between dynamic objects, we propose a 3D alignment loss that effectively accounts for the occlusions and object motions.
@@ 2 Related Work
Figure 2: Cross-attention visualization of DUSt3R wang2024dust3r on static vs. dynamic scenes. We show source-image attention maps for a highlighted query point (red) at each layer and averaged across layers. While DUSt3R captures geometric correspondences well in static regions, it fails in dynamic areas due to its rigid-motion, static-frame assumption.
Figure 3: Cross-attention visualization on dynamic scenes: MonST3R zhang2024monst3r vs. Ours. Using the same setup as Figure 2, MonST3R inherits DUSt3R’s static-scene supervision and fails to align moving regions, limiting reconstruction. In contrast, our method consistently matches dynamic frames and produces sharply localized attention.
@@
Classical 3D reconstruction methods typically follow a multi-stage pipeline to recover scene geometry and camera parameters from a set of uncalibrated images. Prominent examples include Structure-from-Motion (SfM) zhang2022structure , and Simultaneous Localization and Mapping (SLAM) davison2007monoslam . SfM incrementally reconstructs sparse 3D points through feature matching and bundle adjustment where SLAM simultaneously estimates camera trajectories and builds sparse or semi-dense maps in real-time. Building upon SfM, ParticleSfM zhao2022particlesfm incorporate particle-based trajectory modeling to track object motions, paving the way for reconstruction in dynamic scenes. Despite all of these approaches showing remarkable performance, they typically require long processing time and resources for scene-specific optimization and often struggle with error accumulations from multi-stage pipelines.
@@
Building on top of estabilished correspondence fields cho2021cats ; cho2022cats++ ; hong2021deep ; hong2022neural ; hong2022cost , various learning-based approaches wang2024vggsfm ; lin2018learning ; gkioxari2019mesh ; sitzmann2019deepvoxels ; hong2024unifying have been proposed to reconstruct static 3D scenes by learning strong 3D priors, representing the scene as point clouds lin2018learning ; guo2020deep , meshes gkioxari2019mesh ; wang2018pixel2mesh , voxels sitzmann2019deepvoxels ; choy20163d and 3DGS hong2024pf3plat . Recently, DUSt3R wang2024dust3r notably provides a unified, feed-forward pipeline for dense stereo matching, geometry estimation, and triangulation by directly regressing structured 3D pointmaps. Although DUSt3R significantly improves reconstruction quality and efficiency by reducing cumulative errors, DUSt3R is inherently designed for static scenes, limiting its effectiveness in scenarios involving dynamic components.
@@
Dynamic scene reconstruction introduces additional complexities due to non-rigid transformations occurring across frames. Similarly to static scene reconstruction, several recent approaches li2024megasam ; zhang2024monst3r ; lu2024align3r ; liang2025zero ; wang2025continuous ; yang2025fast3r have employed learning-based methods to tackle dynamic scene reconstruction. Among these, MonST3R zhang2024monst3r directly fine-tunes DUSt3R using videos consisting of dynamic scenes. Although it has shown competitive performance, MonST3R retains DUSt3R’s per-frame training paradigm and lacks explicit mechanisms for linking corresponding points across frames in dynamic scenes, leading to inconsistent depth estimations that arises from lacking constraints that fail to capture the intricate motion patterns of objects. Our method addresses this limitation by augmenting the existing feed-forward framework with motion-aware training objectives, explicitly enforcing consistent 3D point correspondences over time.
3 Preliminary
Given a pair of input images $I^{1},I^{2}\in\mathbb{R}^{W\times H\times 3}$, DUSt3R wang2024dust3r predicts a pair of 3D pointmaps $X^{1,1},X^{2,1}\in\mathbb{R}^{W\times H\times 3}$ for both images, each expressed in the camera coordinate system of $I^{1}$. To train the network in a supervised manner, ground-truth pointmaps for each image are defined in the coordinate space of the first camera. Specifically, given the camera intrinsics matrix $K\in\mathbb{R}^{3\times 3}$, world-to-camera pose matrices $P^{n},P^{m}\in\mathbb{R}^{4\times 4}$ for images $n$ and $m$, and a ground-truth depth map $D\in\mathbb{R}^{W\times H}$, the ground-truth pointmap is computed as $\bar{X}^{n,m}=P^{m}(P^{n})^{-1}h(K^{-1}D)$, where $h:(x,y,z)\mapsto(x,y,z,1)$ represents the transformation to homogeneous coordinates.
@@ 3 Preliminary
where $v\in\{1,2\}$ is the input views and $i\in\mathcal{D}^{v}$ denotes valid pixel positions. The scaling factors $z=\text{norm}(X^{1,1},X^{2,1})$ and $\bar{z}=\text{norm}(\bar{X}^{1,1},\bar{X}^{2,1})$ are computed using the normalization function:
@@ 4 Methodology
Figure 4: Construction of alignment loss defined at static and dynamic regions. We introduce a pipeline for constructing a static-dynamic aligned pointmap (SDAP) that explicitly handles occlusions in dynamic regions. First, we compute and refine optical flow via cycle consistency checks and derive a dynamic mask $M_{\text{dyn}}$. To align image $I^{2}$ with image $I^{1}$, we then: 1) warp static pixels using the known camera poses, and 2) warp dynamic pixels using the optical flow. By combining these two warps, we finally obtain SDAP that registers every corresponding 2D pixel into 3D space.
@@ 4.1 Motivation and overview
As shown in Figure 1, DUSt3R wang2024dust3r predicts 3D pointmaps in static regions with high accuracy, leveraging precise stereo correspondences for robust reconstruction. Inspired by ZeroCo’s finding that DUSt3R’s cross-attention inherently encodes geometric correspondences an2024cross , we visualize attention maps for both static and dynamic scenes in Figure 2. While DUSt3R excels in static areas, it fails to produce localized attention scores in dynamic regions, leading to noisy pointmap predictions as visualized in Figure 1. To address this, MonST3R zhang2024monst3r augment DUSt3R’s training data with dynamic sequences, but they still neglect explicit object-to-object correspondences. This omission prevents dynamic objects from serving as anchors that could strengthen the spatial structure and improve depth estimates of neighboring static regions, thereby limiting overall reconstruction quality. Figure 3 highlights these shortcomings, showing how noisy or missing correspondences degrade accuracy.
To overcome these limitations, we propose Static–Dynamic Aligned Pointmaps (SDAP). Unlike static-only aligned pointmaps, which align 3D points solely with rigid transformations, SDAP simultaneously aligns both static and dynamic components in the scene, fully encoding each pixel’s spatial and temporal information. In the following sections, we describe our proposed SDAP representation and outline our training procedure.
4.2 Static-Dynamic Aligned Pointmap
While 3D pointmaps excel at encoding static 3D structure wang2024dust3r , they break down on moving objects due to temporal misalignment. To address this, we introduce Static–Dynamic Aligned Pointmaps (SDAP), which jointly enforce spatial consistency and capture pixel motion across time. Our key insight is to represent every pixel in a unified world coordinate system that aligns its position at each timestep, yielding a cohesive reconstruction of dynamic scenes. However, dynamic motion often leads to occlusions, meaning some pixels appear in only one frame. This results in incomplete alignment, which ultimately limits the quality of 3D reconstruction.
Occlusion Masks.
To eliminate occluded regions in our SDAP representation, we first obtain dense 2D correspondences across each image pair using an off-the-shelf optical flow estimator wang2024sea . If ground-truth flows are available in the dataset, we use those instead. In scenarios with large camera baselines, where occlusions become more frequent, we further apply a forward–backward consistency check sundaram2010dense ; yin2018geonet to derive precise occlusion masks. The full procedure is detailed below.:
$p^{\prime}_{2}=p_{1}+\mathbf{f}(p_{1}),\quad p^{\prime}_{1}=p^{\prime}_{2}+\mathbf{b}(p^{\prime}_{2}),\quad{M}_{\mathrm{occ}}=[|p^{\prime}_{1}-p_{1}|>t],$ (4)
@@ 4.2 Static-Dynamic Aligned Pointmap
Dynamic Masks.
Dynamic regions, characterized by moving objects and non-rigid transformations, introduce significant challenges in accurately aligning pointmaps due to inconsistencies between camera-induced motion and object-specific motion. Without explicitly accounting for such dynamic behaviors, the network may attempt to incorrectly fit dynamic regions as if they were static, thus impairing reconstruction accuracy. To address this, we introduce a dynamic mask ${M}_{\mathrm{dyn}}$ that explicitly highlights moving regions, guiding the network toward a more stable learning process. Specifically, the dynamic mask ${M}_{\mathrm{dyn}}$ is computed by comparing optical flow $\mathbf{f}$ with the expected flow induced purely by camera motion $\mathbf{f}_{\text{cam}}$. Given the depth map $D$, intrinsics matrix $K$, relative rotation and translation $R,T$, and pixel coordinates $p$, we define:
$\mathbf{f}_{\text{cam}}=\pi(DKRK^{-1}p+KT)-p,\quad\mathcal{\mathcal{}}{M}_{\mathrm{dyn}}=[\|\mathbf{f}_{\text{cam}}-\mathbf{f}\|>\tau],$ (5)
where $\pi:(x,y,z)\mapsto(x/z,y/z)$ is the projection operation, and $\tau$ serves as the dynamic threshold.
Table 1: Training datasets. All datasets consists of synthetic scenes and provide both camera and depth. We excluded scenes containing incomplete dataset annotations or noisy objects (e.g., smoke). Details regarding each scene are documented in the supplementary material.
Dataset Domain #\# of frames #\# of Scenes Dynamics Dynamic mask Optical flow Ratio
Blinkvision Outdoor (li2024blinkvision, ) Outdoors 6k 23 Realistic 38.75%
Blinkvision Indoor (li2024blinkvision, ) Indoors 6k 24 Realistic 23.75%
PointOdyssey (zheng2023pointodyssey, ) Indoors & Outdoors 200k 131 Realistic 12.5%
TartanAir (wang2020tartanair, ) Indoors & Outdoors 1000k 163 None 12.5%
Spring (mehl2023spring, ) Outdoors 6k 37 Realistic 12.5%
4.3 Objective Function
Given our SDAP representation equipped with occlusion and dynamic masks, we formulate our training objective as confidence-aware 3D regression wang2024dust3r , leveraging these masks to achieve enhanced stability. To effectively account for the distinct characteristics of static and dynamic regions, we introduce two separate objective functions, each explicitly designed to handle the respective scenarios and ensure accurate reconstruction in both cases.